OpenStudy (anonymous):

Find the mean, median, mode, and range of the data set. 12, 18, 14, 17, 12, 18

OpenStudy (lgbasallote):

do you know what "mean" means?

OpenStudy (anonymous):

Mean(Average) 15.166 Median 15.5 Mode 18 Range 6

OpenStudy (anonymous):

all data better increase order ..

OpenStudy (anonymous):

12,12,14,17,18,18 range : maximum - minimum = 18-12 = 6 Mode : most data = 18 mean (average) = the sum of data / more data = (12+12+14+17+18+18)/6 = 91/6 = 15.166 Median : the middle of data = (14+17)/2 = 31/2 = 15.5